OpenStudy (anonymous):

An amendment to the U.S consitution must be ratified by at least 3/4 of the states to become law rite the ratio as a precent An amendment to the U.S consitution must be ratified by at least 3/4 of the states to become law rite the ratio as a precent @Mathematics

OpenStudy (anonymous):

3/4=x/100 4x=3*100 4x=300 x=300/4= 75%

OpenStudy (anonymous):

u cross multiply the first expression and will get the second